# From attention to profit: quantitative trading strategy based on transformer

### v1

@@ Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Former machine learning approaches have struggled to fully capture various market variables, often ignore long-term information and fail to catch up with essential signals that may lead the profit. This paper introduces an enhanced transformer architecture and designs a novel factor based on the model. By transfer learning from sentiment analysis, the proposed model not only exploits its original inherent advantages in capturing long-range dependencies and modelling complex data relationships but is also able to solve tasks with numerical inputs and accurately forecast future returns over a period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019. The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies with lower turnover rates and a more robust half-life period. Notably, the model’s innovative use transformer to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies.
@@
Quantitative finance, Transformer, Stock selection, Market sentiment
@@ 1 Introduction
The goal of stock trading is to optimize the return on the investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when the difference is generated by the fluctuation of stock spice. However, stocks are influenced by a large number of factors, which constitute a complex system and increase the difficulty for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1]. So, many strategies and tools were built, which corresponded to the development of the capital market and quantitative strategies have played important roles among them.
Figure 1: The Quantitative Trading Architecture with Factors
@@
Some traditional quantitative tools, such as Markowitz Portfolio [2] and Capital Asset Pricing Model(CAPM) [3], focuses mainly on the fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. However, depending on the development of computer science, more quantitative methods and tools were introduced. Within these methods, strategy with factors has attracted people’s attention. In 1993, Fama and French [4] introduced their Fama French Three Factor Model (FF3), which is one of the most influential models in quantitative trading. In 2015, Fama and French renewed their model, the Five-Factor Asset Pricing Model (FF5) [5] was introduced. Besides this classical theory, other trading strategies have been published for decades.
@@
Quantitative trading with factors typically manifests in two primary strategies, which are shown in Fg.1. The first approach involves the computation of stock factor values. Based on these calculated values, stocks are ranked to establish a pool. Once this pool is established, assets are held for a predetermined period. Adjustments to the portfolio are then made at specific time intervals, ensuring alignment with evolving market conditions and factor readings. The second method employs a fixed pool of stocks, wherein factors guide the derivation of long/short signals. Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
@@
In recent years, Machine Learning (ML) stands out as an instrumental tool in enhancing trading algorithms and decision-making processes. Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets. Some traditional literature exploits machine learning algorithms based on manual indicators [6], but the hypothesis about the random walk may be less useful in the task of catching non-stationary series. With the rise of different architectures, quantitative researchers found some novel approaches to identifying intricate patterns and factors in trading data, which opened a new field in quantitative trading [7].
@@
Although there are several previous experiments about ML factors in the quantitative finance field, the research still faces two concerns. Firstly, in the field of NLP, it is necessary to convert words in text to word vectors through word embeddings to serve as inputs. However, in the financial dataset, there are categorical data such as industry types, as well as numerical data like price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence [8]. In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
@@
For the second concern, in NLP tasks, most of them can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. To utilize existing outputs, decoders in NLP tasks sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
@@
To address these concerns, we propose an approach that combines time series stock data with a numerical-adapted transformer to establish a factor and put it into back-tests. This approach allows us to use transformer model embedding numerical input directly, which refers to a method similar to sentiment analysis. The structure of the paper is designed as follows. Section 2 discussed the previous quantitative financial works mainly based on machine learning. Section 3 and 4 introduce the establishment of the model. A factor based on the model will be trained and back-tested. For the practical test, we collected data from more than 4,600 stocks in the previous 14 years (from 2010 to 2023) from the data-collect platform. To comprehensively test the ability of the factor, we divided of data by different frequencies (monthly, weekly, and daily). Finally, the result of the back-tests including the comparison between the transformer-factor and others and the insights gained from such comparative analysis will be discussed in the section 4 and 5.
Figure 2: Example for stock prediction
@@ 2 Preliminaries
This section briefly introduces the related works about the paper and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
2.1 Stock data with market sentiment
Stock data can be considered that can influence the future trend of this stock in some cases [9]. Previous works have discussed the correlation between investor sentiment and stock market metrics like turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns [10, 11]. This relationship isn’t uniform but varies with changes in investor sentiment. Research also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors like the turnover ratio, which are seen as reflective of market liquidity and investor behavior [12]. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
2.2 Applications for ML methods in trading
Machine learning exploits a range of algorithms and statistical techniques used for tasks such as regression, clustering, and classification. In past decades, there have been methods that have been significantly increasingly applied in stock due to their ability to process vast amounts of data and make predictions, offering a potential advantage in the financial markets. At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
@@ 2.2.1 Support Vector Machines (SVM)
SVM was widely used in quantitative trading, which can manage high-dimensional spaces and intricacies found in datasets. In 2003, SVM tried to forecast financial time series [13], especially with the lens focused on the South Korean stock market. This work exhibited the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting. Similar to the previous one, [14] showcased SVM’s prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. [15], found SVM’s adaptability and precision in financial tasks like portfolio and quantitative trading.
However, SVM also shows some disadvantages in quantitative trading. In 2009, SVM was underscored for the accuracy of overfitting [16]. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, causing the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptive [17].
@@ 2.2.2 Long Short-Term Memory (LSTM)
LSTM was designed to capture long-term dependencies in sequential data according to their unique architecture [18]. [19] deployed LSTM with stacked autoencoders to forecast financial time series. LSTM shows the ability to accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. In 2018, [20] tested the efficacy of LSTM for price predictions. They observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over some conventional models. LSTM’s capability to extract features from time series was explored by [21]. They combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, [22] harnessed the power of LSTM to unearth multi-frequency trading patterns. Their research underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
2.2.3 GRU
Gated Recurrent Units (GRU) [23] have emerged as a useful method in the quest for accurate financial forecasting. GRU can capture long-term dependencies in time series data and has seen increasing popularity in the world of quantitative trading. [19] introduced a comprehensive deep learning framework for financial time series, which Incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. [24] centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. What’s more, [25] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [26]. The strategy combined CNNs and GRUs, also received advantages among them, creating a formidable solution for stock price prediction.
@@ 2.3.1 Introduction for transformer
Transformer [27], which was introduced in 2017, has received high attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. The transformer-based models like GPT [28] and BERT [29] are the basis of the specific field of application. In the application field, transformer-based models like ChatGPT [30], FinBERT [31, 32] and BloomberGPT [33] show advantages and ability in solving different types of problems in different filed.
@@
Similar to other neutral sequence transduction models, transformer also has an encoder-decoder structure. The model takes input as a sequence of context which could be natural language sentences and generates output with texts depending on the input, like a translation according to the input or a predicted sentence. Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, there are multi-head attention layers are used and the result from each layer will be outputted after going through an additional &\&& normalize layer. The above-mentioned steps covered most of the progress in the vanilla transformer.
2.3.2 Previous transformer-based strategy
Transformer has achieved superior performance on long-term dependencies modeling. In the quantitative trading domain, some previous researchers tried to test the potential of transformer. [34] firstly tried to exploit Transformer on trading sequences to classify stock price movements.
@@ 2.3.2 Previous transformer-based strategy
Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. [35] devoted to optimizing the efficiency of time complexity and memory usage of Transformer on extremely long time series by informer. [36] leveraged the advantages of CNNs and Transformers to model short-term and long-term dependencies in financial time series. They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents.
@@
Similar to previous work, [1] innovated the structure of a Transformer, which is Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. Finally, they obtained more than 10% of annual return on average.
@@
However, there are some disadvantages of transformer. [37] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, [38] claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
@@ 3 Methodology
This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, model enhancement and output processing.
@@ 3.1 Framework overview
Fig. 3 illustrates the overview of the structure. There are three major parts:
@@
(ii) Embedding the training dataset and training them by transformer which replaces the word-embedding layer to linear layer and deletes the Mask(opt.) layer;

### v2

# Quantformer: from attention to profit with a quantitative transformer trading strategy

@@ Abstract
In traditional quantitative trading practice, navigating the complicated and dynamic financial market presents a persistent challenge. Fully capturing various market variables, including long-term information, as well as essential signals that may lead to profit remains a difficult task for learning algorithms. In order to tackle this challenge, this paper introduces quantformer, an enhanced neural network architecture based on transformers, to build investment factors. By transfer learning from sentiment analysis, quantformer not only exploits its original inherent advantages in capturing long-range dependencies and modeling complex data relationships, but is also able to solve tasks with numerical inputs and accurately forecast future returns over a given period. This work collects more than 5,000,000 rolling data of 4,601 stocks in the Chinese capital market from 2010 to 2019. The results of this study demonstrated the model’s superior performance in predicting stock trends compared with other 100 factor-based quantitative strategies. Notably, the model’s innovative use of transformer-liked model to establish factors, in conjunction with market sentiment information, has been shown to enhance the accuracy of trading signals significantly, thereby offering promising implications for the future of quantitative trading strategies. The implementation details and code is available on https://github.com/QuantFormer.
@@
quantformer, transformer, quantitative finance, stock selection, market sentiment
@@ 1 Introduction
The goal of stock trading is to optimize the return on investment in the capital market according to the process of buying or selling one or more companies’ shares. Traders obtain profit when a positive difference is generated by the fluctuation of stock price. However, stocks are influenced by a large number of factors, which constitute a complex system and make it difficult for people to make a profit. The assessment of a stock’s evolving trend is inherently challenging due to the highly volatile and interconnected nature of the market, which sets it apart from typical time series modeling [1]. As a result, many strategies and tools have been built, in parallel with the development of capital markets, and quantitative strategies have been playing an important role among them.
Figure 1: Quantitative trading architecture with factors
@@
Some traditional quantitative tools, such as the Markowitz portfolio theory [2] and the Capital Asset Pricing Model (CAPM) [3], focus mainly on static fundamental analysis. In other words, these strategies aim to make a profit by simple calculation and analysis. Since then, along with the development of computer science, more quantitative methods and tools have been introduced. Within these methods, factor-based strategies have attracted much attention. In 1993, [4] introduced their Fama-French Three Factor Model (FF3), which has become an influential model in quantitative trading. In 2015, Fama and French revised their model with a Five-Factor Asset Pricing Model (FF5) [5]. Besides this classical theory, numerous trading strategies have been published for decades. [6] built a optimal portfolio to catch future investment opportunities (FIO) by multi-factor models.
@@
Quantitative trading with factors typically follows two primary approaches, which are shown in Figure 1. The first approach involves the computation of stock factor values. Based on these calculated values, stocks are ranked to establish a pool. Once this pool is established, assets are held for a predetermined period. Adjustments to the portfolio are then made at specific time intervals, ensuring alignment with evolving market conditions and factor readings. The second method employs a fixed pool of stocks, wherein factors guide the derivation of long/short signals. Traders can execute corresponding actions when they receive the signals from factors, allowing for a dynamic response to market fluctuations based on factor insights.
@@
In recent years, Machine Learning (ML) has become an instrumental tool for enhancing trading algorithms and decision-making processes. Its core principle, enabling systems to learn from and make decisions based on data, lends itself particularly well to the vast and dynamic landscapes of stock markets. [7] exploits machine learning algorithms based on manual indicators, but the underlying random walk hypothesis may hamper the task of understanding inherently non-stationary series. With the rise of different architectures, [8] finds some novel approaches to catch the sentiment of the market in trading data, which opened a new field in quantitative trading.
@@
Although there exist several previous experiments with ML factors that attempted to fetch market sentiment in the quantitative finance field, this research still faces two difficulties. Firstly, in the field of sentiment analysis, which is a branch of Natural Language Processing (NLP, a field of computer science that aims to understand, interpret, and generate human language), models are used to convert words in text to word vectors through word embeddings to serve as inputs. However, financial datasets contain categorical data instead of words such as industry types, as well as quantitative data such as price fluctuation, turnover rate, and financial indicators. If the input comprises only categorical data, the time series can be treated as a sentence [9]. In most cases, the input will involve numerical data, which cannot be transformed via word embeddings.
@@
Secondly, most NLP tasks can be transformed into sequence-to-sequence (seq2seq) problems, such as in machine translation, dialogue systems, and speech recognition. As an example, the transformer architecture is based on the seq2seq architecture [10]. To utilize existing outputs, decoders in transformer sequentially output samples and use masking operations to handle input sequences during training. However, in stock prediction, where the aim is often to accurately forecast future returns over a period, the transformer model is rarely used for such tasks.
@@
To address these problems, we propose quantformer, which is a modified transformer architecture adapted to quantitative data, and used as an investment factor. Quantformer is able to input numerical data directly, which refers to a method similar to sentiment analysis. The paper is structured as follows. Section 2 discussed the previous quantitative financial works mainly based on machine learning. Sections 3 and 4 introduce quantformers. A factor based on quantformer will be trained and back-tested. For the practical backtest, we collected data from more than 4,600 stocks in the past 14 years (from 2010 to 2023) from the data-collect platform. To comprehensively test the ability of the factor, we divided the data by different frequencies (Section 4.2) and trained under different training scales (Section 5.3). Finally, the result of the back-tests including the comparison between the quantformer-factor and other 100 factors as well as the insights gained from such comparative analysis will be discussed in Sections 4 and 5.
Figure 2: Example of stock prediction
@@ 2 Preliminaries
This section briefly introduces the related works about stock prediction with market sentiment and the development of quantitative financial trading methods, especially based on Machine Learning (ML) methods.
2.1 Stock data and market sentiment
It is known that, in some cases, stock data can be deemed to influence the future trend of a stock price [11]. [12, 13] have discussed the correlation between investor sentiment and stock market metrics such as turnover rate, indicating that investor feelings, both individual and institutional, significantly influence market behavior, affecting aspects such as volatility and returns. This relationship is not uniform but varies with changes in investor sentiment. It also highlights that investor psychology, especially during periods of high market stress can drastically impact market behavior and investor decisions. This link between sentiment and market performance is further affirmed by studies focusing on factors such as the turnover ratio, which are seen as reflective of market liquidity and investor behavior [14]. [15] built a new factor based on sentiment analysis with the average of trading signals from technical trading strategies to benchmark stocks of the S&P 500 index and DJIA. The sentiment factor shows correlation with stock returns. These findings underscore the mutual relationship between investor sentiment and market performance, demonstrating how psychological factors can drive market dynamics.
2.2 Applications of ML methods for trading
Machine learning exploits a range of algorithms and statistical techniques used for tasks such as regression, clustering, and classification. In past decades, there have been methods that have been increasingly applied to stock data due to their ability to process vast amounts of data and make predictions based on them, offering a potential advantage in the financial markets. At the same time, the finance sector, particularly quantitative trading, has started to be aware of the potential of deep learning models for predicting stock movements, portfolio optimization, and risks.
@@ 2.2.1 Support Vector Machines (SVM)
SVM has been used in quantitative trading, as it can manage high-dimensional spaces and intricacies found in datasets. [16] used SVMs to forecast financial time series in 2015, with a focus on the South Korean stock market, exhibiting the robustness of SVM against traditional statistical methods and some other computational techniques, setting a precedent for its adoption in financial forecasting. Similarly, [17] showcased SVMs’ prowess in anticipating the movement direction of Japanese stock markets, especially when determining bullish or bearish market sentiments. [18] assessed SVMs’ adaptability and precision in financial tasks such as portfolio and quantitative trading.
However, SVMs can also show some disadvantages in quantitative trading. In [19], SVM was underscored for the accuracy of overfitting. When applied to financial datasets, the model’s propensity to fit too closely to the training data might lead to reduced generalization capabilities, making the strategy ineffective on unseen data. In high-frequency trading, though the model displayed competency in specific scenarios, researchers pinpointed potential limitations when grappling with certain volatile market dynamics, suggesting the need for adaptations [20].
@@ 2.2.2 Long Short-Term Memory (LSTM)
LSTM is a neural network architecture designed to capture long-term dependencies in sequential data [21]. LSTM neural networks can accentuate and effectively model volatile financial markets, showcasing significant improvements over traditional time series models. [22] combined LSTM with linear Kalman filter and set experiments on different volatility stocks to build strategies. [23] deployed LSTM with stacked autoencoders to forecast financial time series. [24] tested the efficacy of LSTM for price predictions they observed LSTMs’ prowess in noisy data and their superiority in prediction accuracy over other conventional models. [25] combined LSTM with other architectures and found that these can enhance the feature extraction process, thereby improving prediction accuracy. For predicting stock prices, [26] harnessed the power of LSTM to unearth multi-frequency trading patterns. Their method underlined the adaptability of LSTMs in modeling complex situations inherent in stock prices.
2.2.3 Gated Recurrent Units (GRUs)
GRUs [27] are another neural network architecture which has emerged as a method in the quest for accurate financial forecasting. GRUs can capture long-term dependencies in time series data and have seen increasing popularity in the world of quantitative trading. [23] introduced a comprehensive deep-learning framework for financial time series, which incorporated GRUs alongside LSTM and stacked autoencoders. Their model, when applied to diverse datasets, performs better than traditional time series models, accentuating the adaptability and robustness of GRUs. [28] centralized on LSTM networks, provided comparative insights into GRUs. However, in their analysis, LSTMs showed a slight edge in prediction accuracy, indicating that while GRUs are powerful, selecting between them and LSTMs may boil down to specific use cases and computational constraints. Beyond that, [29] presented a model fusing GRUs with Convolutional Neural Networks (CNN) [30]. The strategy to combine CNNs and GRUs also exhibits advantages for stock price prediction. [31] introduced the Attention-GRU model which combined attention and GRU to establish a new factor based on CVaR [32, 33] portfolio. The Attention-GRU model fitted market return, stocks return from 28 Dow Jones Industrial Average index (DJIA) stocks, which achieved better performance than other models in 8 metrics such as annual return, standard deviation and information ratio.
@@ 2.3 Transformer in quantitative trading
Transformer [10], which was introduced in 2017, has received considerable attention in the NLP field. According to its special encoder and decoder stacks with self-attention blocks, the transformer shows its advantage in stability, speed, and long-term memory compared with other traditional models. The transformer-based models such as GPT [34] and BERT [35] are the basis of the specific field of application. In the application field, transformer-based models such as ChatGPT [36], FinBERT [37, 38] and BloomberGPT [39] show advantages and ability in solving different types of problems in different fields. [40] utilized transformer to to predict the stock market index (including CSI 300, S&P 500, Hang Seng Index, and Nikkei 225) and concluded that the model could better catch the rules of stock market dynamics.
@@
Similar to other neural sequence transduction models, transformer also has an encoder-decoder structure. The model takes as input a sequence of context which could be natural language sentences and generates output with texts depending on the input, similar to a translation according to the input or a predicted sentence. Transformer using stacks which are called self-attention and point-wise, which connect layers in the architecture for encoder and decoder. For these encoder and decoder layers, the self-attention method plays an important role in matrix calculation. The linear and softmax functions are used in the dot-product attention module. In the vanilla transformer model, multi-head attention layers are used and the result from each layer will be outputted after going through an additional normalization layer. The above-mentioned steps cover most of the progress in the vanilla transformer.
@@
In the quantitative trading domain, some previous researchers tried to test the potential of transformer. [41] firstly tried to exploit transformer on trading sequences to classify stock price movements. Besides using a traditional transformer, other innovative methods of transformer are worth mentioning. [42] devoted to optimizing the efficiency of time complexity and memory usage of transformer on extremely long time series by informer. [43] takes the advantage of CNNs and transformers to model short-term and long-term dependencies in financial time series. They showed the advantage of the proposed approach on intraday stock price prediction of S&P 500 constituents. [44] proposed a new ConvLSTM model by combining convolutional transformer and LSTM. The model was trained by stocks from S&P 500 from 2004-2021 and performs better than other models such as VaR and ARIMA.
@@
Similar to previous work, [1] innovated the structure of a transformer, in the form of Adaptive Long-Short Pattern Transformer (ALSP-TF). Their model is structurally innovated for hierarchical representation and interaction of stock price series at different context scales. With the help of a learnable function, they make the self-attention aware of the weighted time intervals between patterns, to adaptively adjust their dependencies beyond similarity matching. In the end, they obtained more than 10% of annual return on average.
@@
However, transformer has some disadvantages. [45] mentioned that the global self-attention module focuses on point-wise token similarities without contextual insights. As fluctuations of stocks are conditioned on composite signals over manifold periods, lacking pattern-wise interaction hinders the adequate discrimination of stock tendency and is susceptible to noise points. On the other hand, [46] claimed that the basic query-key matching paradigm is position agnostic. Although position embedding is inserted into the sequential inputs, it may not be optimal because of the inability to reveal precise distances.
@@ 3 Methodology
This section introduces the framework of the work and the steps of the establishment of the model, including the data processing, quantformer construction and prediction.
@@ 3.1 Framework overview
Figure 3 gives an overview of the structure. There are three major parts:
@@
(ii) Embedding the training dataset and training them by a quantformer with word-embedding layer replaced by linear layer and with no Mask(opt.) layer;